OpenStudy (lgbasallote):

A projectile of mass 0.50 Kg is fired with an initial speed of 10 m.s at an angle of \(60^o\) above the horizontal. The potential energy (relative to the ground level) of the projectile at its highest point is...?

OpenStudy (lgbasallote):

do i use \[U_{GRAV} = mgy\]

OpenStudy (anonymous):

Direction : put vertical component of speed to 0 leave horizontal as it is , This the speed at the Apex of the parabola. Calculate the absolute reduction of the E-kinetic this is your potential energy

OpenStudy (anonymous):

Calculate maximum height reached.. Vertical velocity = 10 sin 60 A = -10 calculate from v^2-u^2 = 2as.. multiply height by M*g

OpenStudy (anonymous):

NO NEED to calculate the height directly -ust as I have told you use the reduction i E-kin

OpenStudy (anonymous):

final velocity is zero initial is 10 sin60

OpenStudy (anonymous):

lol.. whatever :P

OpenStudy (lgbasallote):

wait..apex parabola? what's that?

OpenStudy (lgbasallote):

and absolute reduction?

OpenStudy (anonymous):

lmfau - You probably don't realize that one does not need to know the top height to know the max potential enrgy, do ya me LOL..

OpenStudy (lgbasallote):

so much fancy terms

OpenStudy (anonymous):

My remark addresse to vikrang

OpenStudy (anonymous):

Again lgb listen here:

OpenStudy (anonymous):

you can do it either way :S

OpenStudy (anonymous):

Calculate the initial Vertical velocity

OpenStudy (anonymous):

Then lgb, calculate the m* Vvertical^2 /2

OpenStudy (lgbasallote):

initial vertical velocity is 0 right?

OpenStudy (lgbasallote):

oh wait...that's final vertical velocity

OpenStudy (anonymous):

Thsi is Your solution for maximal potential energy

OpenStudy (lgbasallote):

at max height v = 0 right?

OpenStudy (anonymous):

\[V_{y, init} = V*Sin \alpha \]

OpenStudy (anonymous):

And of course at Top V_y = 0 (NOT V though - the body keeps flying)

OpenStudy (lgbasallote):

is that initial V_y = V sin \(\alpha\)

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

so how do i know what V is?

OpenStudy (lgbasallote):

oh 10 m/s

OpenStudy (lgbasallote):

yes. that makes sense

OpenStudy (anonymous):

\[V_{apex}\neq0\]

OpenStudy (lgbasallote):

\[V_{y_i} = 10 \sin (60) \implies 5\sqrt 3\] right?

OpenStudy (lgbasallote):

okay...what's V_apex now?

OpenStudy (anonymous):

The total speed at the top (Apex is Top but latex has Top as a command lol...)

OpenStudy (lgbasallote):

why isnt the v at top 0?

OpenStudy (anonymous):

It should have been simple if you have had calculated height :/

OpenStudy (anonymous):

THIS IS SOMETHING CENTRAL YOU MUST READ IN BOOKS. SHORT ANSWER IS: V-Horizontal is CONSTANT during the whole process. Of course ubder the idealization of no-air-friction

OpenStudy (anonymous):

SO LGB : V_Horizontal is CONSTANT, and in your data it is NONZERO - so V_total is non-zero at the top

OpenStudy (lgbasallote):

lol i agree @vikrantg4 =)) these shortcuts make me confuse

OpenStudy (anonymous):

I repeat\[V_{x} = CONSTANT\]

OpenStudy (lgbasallote):

but at max height it will stop... so velocity there should be 0

OpenStudy (anonymous):

and at the top \[V = V_{x}\]

OpenStudy (anonymous):

Nah.. It stops only going up.. It continues to travel in Horizontal direction

OpenStudy (lgbasallote):

horizontal?

OpenStudy (lgbasallote):

but it's fired at 60 degree angle

OpenStudy (anonymous):

This shows you the projectile DOES NOT STOP AT THE TOP - IT FLIES at that instant of course - HORIZONTALLY

OpenStudy (anonymous):

omg.. at TOP it stops travelling vertically and continues in horizontal direction

OpenStudy (lgbasallote):



OpenStudy (anonymous):

You know resolution of vectors ? eh .. :p

OpenStudy (lgbasallote):

no?

OpenStudy (anonymous):

Lgb : The HOrizontal component of velocity\[V_{x}\] stays OCNSTANT and it is NOT ZERO !

OpenStudy (lgbasallote):

i know the horizontal velocity is constant..what im asking is V_y

OpenStudy (lgbasallote):

V_y should be 0 at max height

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Your statement "the projectile stops" is wrong.

OpenStudy (lgbasallote):

that's what i was aksing...

OpenStudy (anonymous):

V_y at grtound is 10 sin60

OpenStudy (anonymous):

just put conservation of energy and calculate

OpenStudy (anonymous):

It keeps flying with \[V_{x}\]

OpenStudy (lgbasallote):

....so this doesnt use mgy?

OpenStudy (anonymous):

You have TWO methods available A. As I solved for you - just put the difference in kinetic energies and you have the answer B. What Vikrang4 suggets - find the top height and , yes, plug it into m*g*y

OpenStudy (anonymous):

See my detailed suggestion at the beginning of the exchange

OpenStudy (lgbasallote):

to get V i have to do \(\sqrt{V_x^2 + V_y^2}\) first right?

OpenStudy (anonymous):

NOOO

OpenStudy (lgbasallote):

no?

OpenStudy (lgbasallote):

oh wait..why did i do that

OpenStudy (anonymous):

At the top only\[V_{x} \] survuves

OpenStudy (lgbasallote):

nevermind what i said lol

OpenStudy (anonymous):

The \[V_{y} = 0\] at the top point

OpenStudy (lgbasallote):

one more question... \[\huge \Delta U = \Delta K\]??

OpenStudy (anonymous):

Yes but with a small correction - a MINUS sign